OpenStudy (anonymous):

An airplane traveling at half the speed of sound (v = 172 m/s) emits a sound of frequency 5.00 kHz. At what frequency does a stationary listener hear the sound as the plane approaches,and after it passes?

OpenStudy (aaronq):

use these equations http://hyperphysics.phy-astr.gsu.edu/hbase/sound/dopp.html

OpenStudy (anonymous):

thank you :)

OpenStudy (aaronq):

no problem!

OpenStudy (shamim):

I wanna help u

OpenStudy (shamim):

First i wanna tell u abt the frequency which getting listener during approach

OpenStudy (shamim):

Required equation is f'=(v/(v-vs))f

OpenStudy (shamim):

f'=?

OpenStudy (shamim):

Given v=2*172m/s vs=172m/s f=5000hz

OpenStudy (shamim):

Feel free to ask for more clarification!!